OpenStudy (anonymous):

The following function defines a recursive sequence: f(0) = -2 f(1) = 8 f(n) = -4•f(n -1) - 3•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function?

OpenStudy (anonymous):

-2, 8, -26, -80, … -2, 8, -26, 80, … -2, 12, -44, 180, … -2, -12, -44, -180, …

OpenStudy (jango_in_dtown):

see here f(2)=-4.f(1)-3.f(0)=-4.8-3.(-2)=-32+6=-26

OpenStudy (jango_in_dtown):

and f(3)=-4.f(2)-3.f(1)=-4.(-26)-3.8=80

OpenStudy (jango_in_dtown):

option b is correct